Operation Summary for January 12, 2011 - January 14, 2011 |
|
Wednesday Jan. 12 The day shift began with the Tevatron (TeV) in shot setup, with Operations monitoring the Recycler stash (168.6E10), with the Antiproton Source (Pbar) stacking (30.9E10), with the Main Injector (MI), Linac, and Booster operating normally, with the NuMI and MiniBooNE experiments taking beam, with the Meson FTBF T-978 (CALICE) experiment off due to shot setup, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) waiting for the next store. During the day shift, Operations established store 8416 at 8:40 AM with an average initial luminosity (AIL) of 349.48E30. NuMI personnel conducted target and horn scan studies. During the evening shift, Operations disabled beam to T-978 at 6:08 PM. Operations terminated store 8416 at 11:57 PM. Thursday Jan. 13 The midnight shift began with Operations monitoring the stash (355.4E10), with Pbar stacking (19.9E10), and with NuMI and MiniBooNE taking beam. During the midnight shift, Operations put the TEV into HEP shot setup at 12:12 AM and established store 8417 at 1:15 AM with an AIL of 331.01E30. Operations rebooted MECAR, which caused a few minutes of downtime. Operations established beam to T-978 at 4:25 AM. During the day shift, Operations reported at 10:21 AM that store 8417 aborted due to a TEV RF station (TRF4) tripping off. RF technicians replaced a faulty flow turbine on TRF4 and began conditioning TRF8. Operations put the TEV into shot setup at 12:15 PM and established store 8418 at 1:44 PM with an AIL of 275.63E30. FESS personnel worked on the CUB processed chilled water system. During the evening shift, Operations disabled beam to T-978 at 6 PM. Operations reported several short downtimes due to problems with Linac and Booster RF systems. Friday Jan. 14 The midnight shift began with Operations monitoring store 8418 (80.93E30) and the stash (269.8E10), with Pbar stacking (7.5E10), and with NuMI and MiniBooNE taking beam. During the midnight shift, Operations established beam to T-978 at 4:16 AM. Operations terminated store 8418 at 5:06 AM, put the TEV into shot setup, and established store 8419 at 6:28 AM with an AIL of 315.25E30. Operations reported some minor down times caused by Linac and MI RF trips. The Plans for Friday and the Weekend The plans for today and the weekend are to stack, stash, and store. |
|
More Information Comments and Suggestions |
| last modified 01/14/2011 email Fermilab |
| Security, Privacy, Legal | |